# ZenSVI: An Open-Source Software for the Integrated Acquisition, Processing and Analysis of Street View Imagery Towards Scalable Urban Science

### v1

@@ 3.2 Metadata Analysis
Metadata is important in facilitating exploratory data analyses and for adding more context to a study. This is particularly important for crowdsourced SVI because they are usually captured in diverse ways, unlike those from commercial platforms. Such inconsistency makes it difficult for researchers to know if a specific image is suitable for their research when they acquire numerous SVI (e.g., millions at the city scale). For example, if one is interested in nighttime walkability, it is unsuitable to use SVI from daytime or SVI captured on highways without sidewalks (Hou et al., 2024).
As exemplified in Figure 2, this sub-package for metadata analysis builds upon the basic metadata provided by Mapillary (such as capture time, camera parameters, and spatial coordinates) and further augments it with additional analytical capabilities. The package provides solutions for comprehensive metadata analysis at three levels: 1) image-level, 2) grid-level, and 3) street-level analyses.
@@
For the street- and grid-level analyses, we match each SVI point with the nearest street and the corresponding grid cell to compute indicators at the aggregated level (see Table 4), which are commonly used in use cases (e.g., average sky view factor at the district scale) and could also be useful for data quality analyses such as spatial and temporal coverage (Hou and Biljecki, 2022). This sub-package also contributes to the standardization of data properties by enabling researchers to explicitly query SVI based on both the original Mapillary metadata and our augmented indicators, allowing them to fetch only images that are suitable for their needs.
@@ 3.3 Transformation
Transformation of SVI is important because different image projections and data formats are suitable for different purposes (Biljecki et al., 2023). Our sub-package for image transformation offers options to transform panoramic imagery into two further view types: ‘perspective’ and ‘fisheye’. An overview and definitions of different image projections are provided in Figure 3.
@@
Figure 4: Example of depth estimation for a Mapillary SVI using the Depth Anything model and the resulting point cloud generation.
For point cloud generation, we implement the methodology of Cavallo (2015), which combines depth estimates with RGB color information. The process projects pixels into spherical coordinates with depth-determined radii and then converts them to Cartesian coordinates for three-dimensional positioning. As demonstrated in Figure 4, this workflow transforms panoramic SVI into detailed point clouds, enabling researchers to conduct advanced spatial analyses such as sidewalk width measurement and tree volume estimation (Tu et al., 2024; Orozco Carpio et al., 2024). The visualization shows both top and side views, revealing the scene’s outline, direction, and relative distances between street objects.
@@ 3.4 Computer vision
Our package provides both semantic and panoptic segmentation capabilities. While semantic segmentation classifies pixels into categories, panoptic segmentation additionally identifies unique object instances, enabling tasks like counting individual trees or cars — a crucial feature for scalable urban analysis (Sánchez and Labib, 2024). We employ the state-of-the-art Mask2Former model (Cheng et al., 2022), pre-trained on two key urban datasets. The model achieves 84.5% mIoU for semantic and 66.6% for panoptic segmentation on Cityscapes, and 64.7% mIoU for semantic and 45.5% for panoptic segmentation on Mapillary Vistas.
@@
As demonstrated in Figure 6, the segmentation results clearly identify and color-code various street features including buildings, roads, sidewalks, and vegetation. The package outputs pixel ratios for all labels to a CSV file, facilitating quantitative analysis (Ibrahim et al., 2020). This functionality enables measurement of street elements across different locations and time periods (Ye et al., 2019; Ramírez et al., 2021; Ma et al., 2021), supporting applications such as green view index calculation — a measure of visible greenery at eye-level (Li et al., 2015; Yang et al., 2009).
Our package provides several pre-trained image classification models, with particular emphasis on street-level perception analysis. A key model is trained on the PlacePulse2.0 dataset (Dubey et al., 2016), which comprises 110,988 images from 56 cities across 28 countries. This dataset captures the human perception of streetscapes along six dimensions: depressing, boring, beautiful, safe, lively, and wealthy, based on volunteer-provided labels. Following the methodology of Kang et al. (2023), we implement a ResNet50 architecture where perceptual scores are categorized into five classes. The models achieved an average top-2 accuracy of 66.03% across all six perceptual dimensions. When processing input images, the package generates a five-dimensional probability vector for each perceptual category, which is then converted into final dimensional scores ranging from 0 to 10. An existing model used in Global Streetscapes (Hou et al., 2024) was also added for benchmarking purposes. This model implements a Vision Transformer (ViT) and also generates perception scores, ranging from 0 to 10, with an average accuracy ranging from 61.6% to 77.1% for all six perceptual dimensions.
@@
Our package also implements scene classification using the Places365 dataset (Zhou et al., 2017), which encompasses 365 scene categories (e.g., residential neighborhood, building facade, field road) and 102 scene attributes (e.g., biking, asphalt, grass), and the ResNet model developed by Zhou et al. (2017), which achieves 85.08% precision in top-5 scene category classification. This model has been extensively utilized in various studies for environmental feature extraction (Luo et al., 2023; Yang and Zhang, 2024) and specific scene filtering (Chen et al., 2020a; de Oliveira et al., 2020; Seresinhe et al., 2017).
As demonstrated in Figure 8, our implementation is capable of analyzing street view images across diverse settings (highway, residential neighborhood, slum, campus, and farm), generating both probability distributions over scene categories and corresponding heatmaps. These heatmaps highlight image regions that most strongly influence the classification (Zhang et al., 2018). For example, highway classifications emphasize road infrastructure features, while residential neighborhood analyses focus on foliage and building elements. This functionality enables researchers to conduct detailed environmental analyses through automated scene recognition (Chen et al., 2020b; Yang and Zhang, 2024).
@@
In our computer vision sub-package, we also integrate image classification models developed by Hou et al. (2024), trained on the manually labeled subset of the NUS Global Streetscapes dataset comprising more than 10,000 street view images from over 400 cities across all continents. These images are sourced from Mapillary and KartaView platforms. Our sub-package implements eight distinct classification models that analyze various image attributes:
@@
As illustrated in Figure 9, these models enable comprehensive filtering of SVI based on specific research requirements. For instance, researchers studying nighttime pedestrian experiences can isolate high-quality images taken from walking surfaces during nighttime conditions (Hou et al., 2024). This systematic filtering capability helps mitigate potential biases in street view image analysis.
@@
Our sub-package implements image embedding capabilities, leveraging various pre-trained models including ResNet, VGG, DenseNet, AlexNet, and EfficientNet (He et al., 2016; Simonyan and Zisserman, 2014; Huang et al., 2017; Krizhevsky et al., 2012; Tan and Le, 2019). These models transform street view images into multi-dimensional vector representations that capture their essential features. The embedding process extracts high-level features from the models’ final layers. For instance, ResNet generates a 2048-dimensional vector from its average pooling layer, as illustrated in Figure 10. To quantify image similarities, we implement cosine distance metrics, defined as:
@@ 3.5 Visualization
The visualization sub-package, built with Matplotlib (Hunter, 2007), provides four high-level plotting functions to streamline the presentation of SVI analysis for quick initial data exploration. The package supports histogram generation for discrete data visualization, such as object detection counts, and kernel density estimation plots for continuous variables, such as semantic segmentation ratios. Users can create image grids to display SVI samples, either randomly selected or sorted by specified variables.
@@ 4 Examples and case study
In this section, we conduct a comprehensive case study to showcase examples of how our package can be used to help researchers in one of the most typical studies by exhibiting each functionality in our package. We use Singapore as the study site and aim to invetigate the daytime urban characteristics by conducting k-means clustering based on the features extracted from SVI obtained from Mapillary. This analysis was chosen because this type of analysis with SVI has been implemented for various purposes by researchers (Fan and Biljecki, 2024; Gao et al., 2024; Zhang et al., 2023; Liang et al., 2023; Wang et al., 2024a), and it also includes data engineering processes such as filtering unsuitable imagery. This process was conducted as follows:
@@ 4.1 Download and quality control
We downloaded SVI from Mapillary by simply inputting ‘Singapore’ to the download_svi function, which was then passed to OSMnx to retrieve the boundary of Singapore. This yielded 932,596 images as of June 2024. Then, we applied the following computer vision models to extract features, which we used for the subsequent cleaning process and analysis:
@@ 5.1 Extending ZenSVI
Moreover, the computer vision sub-package can also be extended by adding new deep learning models, such as a zero-shot segmentation model (e.g., the Segment Anything model (Kirillov et al., 2023)) and various generative models, such as text-to-image (Saharia et al., 2022), image captioning (Li et al., 2022a), image-to-sound (Zhuang et al., 2024), image-to-video (Singer et al., 2022), image-to-image (Ito et al., 2024c), and image inpainting (Lugmayr et al., 2022). The visualization sub-package can also be developed further by adding different types of plots based on one’s needs (e.g., alluvial plots for time-series SVI analysis).
@@
5.2 Export of the data, integration, and downstream analyses thanks to ZenSVI
Most of the functions in ZenSVI allow users to save outputs directly to external files. Thus, it is easy to integrate the analysis in our package with other packages. For example, users can download the SVI and perform semantic segmentation to compute the green view index, which can be saved to a CSV file. This design enables users to perform subsequent analysis and visualization in external software. For example, one can conduct a statistical analysis of the index in R and visualize the results in QGIS. Such a bridge facilitates conducting analyses and visualization in their most familiar environment while efficiently conducting the SVI analysis in ZenSVI.
@@ 6 Conclusion
SVI has gained popularity in many research domains for its scalability, spatial resolution, and capability to capture street-level information with high veracity and usability. Despite many applications demonstrated in the literature, the lack of standardized, cohesive, and reproducible software makes it difficult for other researchers to reproduce or expand existing methodologies that rely on such data and develop new ones. To fill this gap, we designed and implemented ZenSVI — an integrated Python package that incorporates a comprehensive set of tools, simplifying the process of SVI analysis and contributing to the acceleration of research relying on this emerging and powerful geospatial dataset. We release this software as open-source and support it with extensive documentation and examples.
@@ Acknowledgments
This research was funded by the Singapore International Graduate Award (SINGA) scholarship provided by the Agency for Science, Technology, and Research (A*STAR), the NUS Graduate Research Scholarship (GRSUR0600042), and the President Graduate Fellowship, all granted by the National University of Singapore (NUS). This research has been supported by Takenaka Corporation. This research has been supported by the Research Activities Fund of City University of Hong Kong. This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant R-295-000-171-133. This research is part of the project Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1. The research was partially conducted at the Future Cities Lab Global at the Singapore-ETH Centre, which was established collaboratively between ETH Zürich and the National Research Foundation Singapore (NRF) under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. We express our gratitude to the members of the NUS Urban Analytics Lab for their valuable discussions and insights. We would like to thank the developers of the open-source software packages that made ZenSVI possible. We also acknowledge the contributors of OpenStreetMap, Mapillary, and KartaView and other platforms for providing valuable open data resources that support street-level imagery research and applications.
@@
During the preparation of this work the author(s) used ChatGPT in order to format tables and proofread. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.

### v2

@@ 3.2 Metadata Analysis
Metadata is important in facilitating exploratory data analyses and for adding more context to each image. This is particularly important for crowdsourced SVI because they are usually captured in diverse ways, unlike those from commercial platforms. Such inconsistency makes it difficult for researchers to know if a specific image is suitable for their research when they acquire numerous SVI (e.g., millions at the city scale). For example, if one is interested in nighttime walkability, it is unsuitable to use SVI from daytime or SVI captured on highways without sidewalks (Hou et al., 2024).
As exemplified in Figure 2, this sub-package for metadata analysis is built upon the basic metadata provided by Mapillary (such as capture time, camera parameters, and spatial coordinates) and further augments it by computing additional features. The package provides solutions for comprehensive metadata analysis at three levels: 1) image-level, 2) grid-level, and 3) street-level analyses.
@@
For the street- and grid-level analyses, we match each SVI point with the nearest street and the corresponding grid cell to compute indicators at the aggregated level (see Table 4), which are commonly used in use cases (e.g., average sky view factor at the district scale) and could also be useful for data quality analyses such as spatial and temporal coverage (Hou and Biljecki, 2022). This sub-package also contributes to the standardization of data properties by enabling researchers to explicitly query SVI based on both the original Mapillary metadata and our augmented indicators, allowing them to fetch only images suitable for their needs.
@@ 3.3 Transformation
Transformation of SVI is also important because different image projections and data formats are required for different purposes  (Biljecki et al., 2023). Our sub-package for image transformation offers options to transform panoramic imagery into two further view types: ‘perspective’ and ‘fisheye’. An overview and definitions of different image projections are provided in Figure 3.
@@
Figure 4: Example of depth estimation for a Mapillary SVI using the Depth Anything model and the resulting point cloud generation. Source of imagery: Mapillary.
For point cloud generation, we implemented the methodology of Cavallo (2015), which combines depth estimates with RGB color information. The process projects pixels into spherical coordinates with depth-determined radii and then converts them to Cartesian coordinates for three-dimensional positioning. As demonstrated in Figure 4, this workflow transforms panoramic SVI into detailed point clouds, enabling researchers to conduct advanced spatial analyses such as sidewalk width measurement and tree volume estimation (Tu et al., 2024; Orozco Carpio et al., 2024). The visualization shows both top and side views, revealing the scene’s outline, direction, and relative distances between street objects.
@@ 3.4 Computer vision
Our package provides both semantic and panoptic segmentation capabilities. While semantic segmentation classifies pixels into categories, panoptic segmentation additionally identifies unique object instances, enabling tasks like counting individual trees or cars — a crucial feature for scalable urban analysis (Sánchez and Labib, 2024). We use the Mask2Former model (Cheng et al., 2022), pre-trained on two key urban datasets. The model achieves 84.5% mIoU for semantic and 66.6% for panoptic segmentation on Cityscapes, 64.7% mIoU for semantic and 45.5% for panoptic segmentation on Mapillary Vistas.
@@
As demonstrated in Figure 6, the segmentation results clearly identify and color-code various street features including buildings, roads, sidewalks, and vegetation. The package outputs pixel ratios for all labels to a CSV/JSON file, facilitating quantitative analysis (Ibrahim et al., 2020). This functionality enables measurement of street elements across different locations and periods (Ye et al., 2019; Ramírez et al., 2021; Ma et al., 2021), supporting applications such as green view index calculation — a measure of visible greenery at eye-level (Li et al., 2015; Yang et al., 2009).
Our package provides several pre-trained image classification models, with particular emphasis on street-level perception analysis. A key model is trained on the Place Pulse 2.0 dataset (Dubey et al., 2016), which comprises 110,988 images from 56 cities across 28 countries. This dataset captures the human perception of streetscapes along six dimensions: depressing, boring, beautiful, safe, lively, and wealthy, based on volunteer-provided labels. Following the methodology of Kang et al. (2023), we implemented a ResNet50 architecture where perceptual scores are categorized into five classes. The models achieved an average top-2 accuracy of 66.03% across all six perceptual dimensions. When processing input images, the package generates a five-dimensional probability vector for each perceptual category, which is then converted into final dimensional scores ranging from 0 to 10. An existing model used in Global Streetscapes (Hou et al., 2024) was also added for benchmarking purposes. This model has a Vision Transformer (ViT) architecture and also generates perception scores, ranging from 0 to 10, with an average accuracy ranging from 61.6% to 77.1% for all six perceptual dimensions.
@@
ZenSVI incorporates a scene classification model pre-trained on the Places365 dataset (Zhou et al., 2017) by Zhou et al. (2017), which achieves 85.08% top-5 precision in scene category classification. This comprehensive dataset contains 365 distinct scene categories (such as residential neighborhoods, building facades, and field roads) and 102 scene attributes (including features like biking areas, asphalt surfaces, and grass coverage). The model’s usefulness has been validated through numerous research applications, from environmental feature extraction (Luo et al., 2023; Yang and Zhang, 2024) to targeted scene filtering for specific urban analyses (Chen et al., 2020a; de Oliveira et al., 2020; Seresinhe et al., 2017).
As demonstrated in Figure 8, our implementation excels at analyzing SVI across diverse urban and rural settings, including highways, residential neighborhoods, slums, campuses, and farms. For each image, the model generates both probability distributions across scene categories and class activation mapping (CAM) heatmaps. These CAM heatmaps visually reveal which regions of the image most significantly contribute to the scene classification (Zhang et al., 2018). For example, the model focuses on road infrastructure and traffic patterns when classifying highways, while for residential neighborhoods, it attends to architectural features, landscaping elements, and the overall spatial organization of buildings and vegetation. This functionality enables researchers to conduct detailed environmental analyses through automated scene recognition (Chen et al., 2020b; Yang and Zhang, 2024).
@@
In our computer vision sub-package, we also integrated image classification models developed by Hou et al. (2024), trained on the manually labeled subset of the NUS Global Streetscapes dataset comprising more than 10,000 images from over 400 cities across all continents. These images were sourced from Mapillary and KartaView platforms. Our sub-package has eight distinct classification models that analyze the following image attributes:
@@
As illustrated in Figure 9, these models enable comprehensive filtering of SVI based on specific research requirements. For instance, researchers studying nighttime pedestrian experiences can isolate high-quality images taken from walking surfaces during nighttime conditions (Hou et al., 2024). This systematic filtering capability helps mitigate potential biases in SVI analysis.
@@
For object detection tasks, we integrate GroundingDINO in ZenSVI. This state-of-the-art model can detect arbitrary objects based on human inputs such as category names or referring expressions (Liu et al., 2023). Trained on several widely used object detection datasets — including COCO, O365, and Open Images — GroundingDINO has also been implemented in urban research, enabling large-scale analyses of built environment features from SVI (Liang et al., 2024; Ramalingam and Kumar, 2025). By specifying category names (e.g., ‘tree’, ‘building’, or multiple objects such as ‘person . car .’), this open-set detector can generate bounding boxes for the identified objects in images. This open-set capability enables the quantification of urban elements and offers flexibility for a wide range of applications.
Our sub-package also offers image embedding capabilities, leveraging various pre-trained models including ResNet, VGG, DenseNet, AlexNet, and EfficientNet (He et al., 2016; Simonyan and Zisserman, 2014; Huang et al., 2017; Krizhevsky et al., 2012; Tan and Le, 2019). These models transform SVI into multi-dimensional vector representations that capture their essential features. The embedding process extracts high-level features from the models’ final layers. For instance, ResNet generates a 2048-dimensional vector from its average pooling layer, as illustrated in Figure 10. To quantify image similarities, we implemented cosine distance metrics, defined as:
@@ 3.5 Visualization
The visualization sub-package, built with Matplotlib (Hunter, 2007), provides four high-level plotting functions to streamline the presentation of SVI analysis for quick initial data exploration. ZenSVI supports histogram generation for discrete data visualization, such as object detection counts, and kernel density estimation plots for continuous variables, such as semantic segmentation ratios. Users can create image grids to display SVI samples, either randomly selected or sorted by specified variables.
@@ 4 Examples and case study
In this section, we conducted a comprehensive case study to showcase examples of how our package can be used to help researchers in one of the most typical studies by exhibiting each functionality in our package. We used Singapore as the study site and aim to investigate the daytime urban characteristics by conducting k-means clustering based on the features extracted from SVI obtained from Mapillary. This analysis was chosen because this type of analysis with SVI has been implemented for various purposes by researchers (Fan and Biljecki, 2024; Gao et al., 2024; Zhang et al., 2023; Liang et al., 2023; Wang et al., 2024a), and it also includes data engineering processes such as filtering unsuitable imagery. This process was conducted as follows:
@@ 4.1 Download and quality control
We downloaded SVI from Mapillary by simply inputting ‘Singapore’ to the download_svi function, which was then passed to OSMnx to retrieve the boundary of Singapore. This yielded 932,596 images as of June 2024. Then, we applied the following computer vision models to extract features used for the subsequent cleaning process and analysis:
@@ 5.1 Extending ZenSVI
The computer vision sub-package can be extended with new models, such as zero-shot segmentation (Kirillov et al., 2023) and various generative models for text-to-image (Saharia et al., 2022), image captioning (Li et al., 2022a), image-to-sound (Zhuang et al., 2024), image-to-video (Singer et al., 2022), image-to-image (Ito et al., 2024c), and inpainting (Lugmayr et al., 2022). The visualization sub-package can also be expanded with additional plot types like alluvial plots for time-series analysis. While leveraging ecosystems like HuggingFace, ZenSVI provides an abstraction layer that simplifies data handling while supporting models from multiple platforms through a consistent interface.
@@
5.2 Import and export of the data, integration, and downstream analyses thanks to ZenSVI
The modularized structure of ZenSVI, with its independent download and other processing sub-packages, allows users to flexibly import and export data that work with external data sources and software tools. For example, users can apply ZenSVI’s transformation and computer vision sub-packages to SVI from other sources not covered in the current version, such as Baidu Street View and Bing Maps. Most of the functions in ZenSVI also allow users to save outputs directly to external files. Thus, it is easy to integrate the analysis in our package with other packages. For example, users can download the SVI and perform semantic segmentation to compute the green view index, which can be saved to a CSV file. This design enables users to perform subsequent analysis and visualization in external software. For example, one can conduct a statistical analysis of the index in R and visualize the results in QGIS. Such a bridge facilitates conducting analyses and visualization in their most familiar environment while efficiently conducting the SVI analysis in ZenSVI.
@@ 6 Conclusion
SVI has gained popularity in many research domains for its scalability, spatial resolution, and capability to capture street-level information with high veracity and usability. Despite many applications demonstrated in the literature, the lack of standardized, cohesive, and reproducible software makes it difficult for other researchers to reproduce or expand existing methodologies that rely on such data and develop new ones. To fill this gap, we designed and implemented ZenSVI — an integrated Python package that incorporates a comprehensive set of tools, simplifying the process of SVI analysis and contributing to the acceleration of research relying on this emerging and powerful geospatial dataset.
@@ Acknowledgments
This research was funded by the Singapore International Graduate Award (SINGA) scholarship provided by the Agency for Science, Technology, and Research (A*STAR), the NUS Graduate Research Scholarship, and the President Graduate Fellowship, all granted by the National University of Singapore (NUS). This research has been supported by Takenaka Corporation. This research has been supported by the Research Activities Fund of City University of Hong Kong. This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant R-295-000-171-133. This research is part of the project Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1. The research was partially conducted at the Future Cities Lab Global at the Singapore-ETH Centre, which was established collaboratively between ETH Zürich and the National Research Foundation Singapore (NRF) under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. We express our gratitude to the members of the NUS Urban Analytics Lab for their valuable discussions and insights. We would like to thank the developers of the open-source software packages that made ZenSVI possible. We also acknowledge the contributors of OpenStreetMap, Mapillary, and KartaView, and other platforms, for providing valuable open data resources that support street-level imagery research and applications.
@@
During the preparation of this work, the author(s) used ChatGPT in order to format tables and proofread. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.